### v1

@@ 3.1. Shelling Order
    Now, let $F^{\prime\prime}\notin M(\Delta_{3}(W_{n}))$, i.e., $W_{n}[F^{\prime\prime c}]$ is connected. We have $i_{1}<j_{1}<\alpha_{j}$ and $22\leq\alpha_{j}\leq n-$ 2. Suppose $j_{1}\nsim i_{1}$ or $j_{1}\nsim\alpha_{j}$. Then $i_{1}\sim\alpha_{j}$ such that $i_{1}=0$ and $\alpha_{j}=n-2$. Therefore, $\alpha_{j}<_{\mathcal{O}}i_{2}$ implies that $i_{2}\in\{1,2,n-1\}$ by definition of $\mathcal{O}$. We get a contradiction to the fact that $W_{n}[F_{i}^{c}]$ is disconnected.
@@
    Thus, $j_{1}\sim i_{1}$ and $j_{1}\sim\alpha_{j}$. Now, since $W_{n}[F_{j}^{c}]$ is disconnected, $j_{2}\nsim\alpha_{j}$, which implies that $j_{2}\geq\alpha_{j}+3$. We have ${F^{\prime}}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Let $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Since $i_{1}<j_{1}<\alpha_{j}$ and $j_{2}\geq\alpha_{j}+3$, $F^{\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, and thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Hence, by taking $F_{r}=F^{\prime}$, ($\ast$3.1) is satisfied.
@@
    We now assume that $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, i.e., $W_{n}[F^{\prime c}]$ is connected. Since $j_{2}\nsim\alpha_{j}$, we get $j_{2}\sim i_{1}$ and $i_{1}\sim\alpha_{j}$. Then $i_{1}<j_{1}<\alpha_{j}<j_{2}$ implies that $i_{1}=\alpha_{j}-2$ and $j_{1}=\alpha_{j}-1$. Also, $j_{2}\sim i_{1}$ implies that either $j_{2}=n-2$ and $i_{1}=0$, or $j_{2}=n-1$ and $i_{1}\in\{0,1\}$. If $i_{1}=0$, then $\alpha_{j}=2$, and thus, $\alpha_{j}<_{\mathcal{O}}i_{2}$ implies that $i_{2}\in\{1,n-2,n-1\}$ by definition of $\mathcal{O}$. This means that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. So, $i_{1}=1$, $j_{1}=2$, $\alpha_{j}=3$ and $j_{2}=n-1$. Further, $W_{n}[F_{i}^{c}]$ is disconnected implies that $i_{2}\nsim\alpha_{j}$, and since $i_{2}\neq j_{2}$, $5<i_{2}<n-1=j_{2}$. It follows that $i_{2}\nsim j_{1}$. We have ${F^{\prime\prime\prime}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Clearly, $W_{n}[F^{\prime\prime\prime c}]\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-1$, $i_{2}\geq\alpha_{j}+3$ and $\alpha_{j}=3\leq m-2$ (as $m\geq 5$), we have $F^{\prime\prime\prime}\notin\mathcal{D}$. Also, $i_{2}<j_{2}$ implies that $F^{\prime\prime\prime}\ll F_{j}$, and thus $F^{\prime\prime\prime}\prec F_{j}$ by Definition 3.4 (i). Hence, $F_{r}=F^{\prime\prime\prime}$ satisfies ($\ast$3.1).
@@
We have ${F_{i}}^{c}=\{\alpha_{i}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$ such that $F_{i}\prec F_{j}$. Then $F_{j}\notin\mathcal{D}$ and $F_{i}\notin\mathcal{T}_{j_{0}}$ implies that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ by Definition 3.4.
Let ${F_{q}}^{c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. Observe that if $F_{q}\in M(\Delta_{3}(W_{n}))$, then $F_{q}\prec F_{j}$ by Proposition 3.5, and hence $F_{r}=F_{q}$ satisfies ($\ast$3.1) by Proposition 3.6.
@@
In this case, we show that $F_{q}\in M(\Delta_{3}(W_{n}))$, and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1).
@@
    We have $j_{1}<j_{2}<\alpha_{j}<\alpha_{i}<m<n-1$. Thus $j_{1}\nsim\alpha_{i}$. Also, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. It follows that $F_{q}\in M(\Delta_{3}(W_{n}))$.
@@
    Since $\alpha_{j}<\alpha_{i}<m$, $\alpha_{j}\leq m-2$. Further, $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}>\alpha_{j}$ implies that $j_{1}\geq m+2$ by definition of $\mathcal{O}$. We have $\alpha_{j}<\alpha_{i}<m<m+2\leq j_{1}<j_{2}$. Thus $j_{1}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$.
@@
    Since $\alpha_{j}<\alpha_{i}<m$, $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}>\alpha_{j}$, we get $j_{2}\geq m+2$. We have $j_{1}<\alpha_{j}<\alpha_{i}<m<m+2\leq j_{2}$. Thus $j_{2}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
In this case, we show that either $F_{q}\in M(\Delta_{3}(W_{n}))$ and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1), or $F_{q}\notin M(\Delta_{3}(W_{n}))$ and there exists some $F_{r}\neq F_{q}$ that satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
    Observe that since $j_{1},j_{2}<\alpha_{j}$, if $\alpha_{j}<4$, then $W_{n}[F_{j}^{c}]$ is connected. So $\alpha_{j}\geq 4$, and hence $\alpha_{i}<2m-\alpha_{j}\leq 2m-4\leq n-3$. We have $j_{1}<j_{2}<\alpha_{j}<m\leq\alpha_{i}<n-3$. Thus $j_{1}\nsim\alpha_{i}$. Since $W_{n}[F_{j}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$.
@@
    Observe that if $j_{1}\nsim\alpha_{i}$ or $j_{1}\nsim j_{2}$, then $\alpha_{i}\geq m$ implies that $j_{2}\nsim\alpha_{i}$, and hence $F_{q}\in M(\Delta_{3}(W_{n}))$. So assume that $j_{1}\sim\alpha_{i}$ and $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{\alpha_{j},j_{2}\}$. We first show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    Since $W_{n}[F_{j}^{c}]$ is disconnected, $j_{1}\sim j_{2}$ implies that $j_{1}\nsim\alpha_{j}$ and $j_{2}\nsim\alpha_{j}$. Therefore, if $\alpha_{i}\nsim\alpha_{j}$ or $\alpha_{i}\nsim j_{2}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $\alpha_{i}\sim\alpha_{j}$. We need to show that $\alpha_{i}\nsim j_{2}$. Since $\alpha_{j}<m$ and $\alpha_{i}\geq m$, $\alpha_{i}\sim\alpha_{j}$ implies that $\alpha_{j}\in\{m-2,m-1\}$. If $\alpha_{j}=m-2$, then $\alpha_{i}=m$ and $j_{2}>j_{1}\geq m+2$ (as $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}>\alpha_{j}$). Thus $\alpha_{i}\nsim j_{2}$.
@@
    Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, from Proposition 3.5, we see that $F^{\prime}\prec F_{j}$. Thus $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
@@
    Suppose $j_{2}\nsim\alpha_{i}$. Now, since $W_{n}[{F_{j}}^{c}]$ is disconnected, we have $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{2}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$; and if $j_{1}\nsim\alpha_{j}$, then $j_{1}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $W_{n}[F_{q}^{c}]$. Thus, we see that $F_{q}\in M(\Delta_{3}(W_{n}))$.
    Now, we assume that $j_{2}\sim\alpha_{i}$. Suppose $j_{1}\sim\alpha_{i}$. Then, $j_{1}<\alpha_{j}<m\leq\alpha_{i}<j_{2}$ implies that either $j_{1}=m-2$, $\alpha_{j}=m-1$, $\alpha_{i}=m$ and $j_{2}\in\{m+1,m+2\}$, or $j_{1}=0$, $\alpha_{i}=n-2$, and $j_{2}=n-1$. In the former case, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{2}=m+2=\alpha_{j}+3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{2})$, a contradiction as $F_{j}\notin\mathcal{D}$. Hence, we have $j_{1}=0$, $\alpha_{i}=n-2$, and $j_{2}=n-1$. Now, $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}<m$ implies that $\alpha_{j}\leq 2$. It follows that $W_{n}[{F_{j}}^{c}]$ is connected, again a contradiction. Therefore, $j_{1}\nsim\alpha_{i}$. Now, if $j_{1}\nsim j_{2}$, then $F_{q}\in M(\Delta_{3}(W_{n}))$.
    So, let $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{j_{1},\alpha_{j}\}$. Since $W_{n}[F_{j}^{c}]$ is disconnected and $j_{1}\sim j_{2}$, we get $j_{1}\nsim\alpha_{j}$. Therefore, $j_{1}\nsim\alpha_{i}$ implies that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. We have $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$. Thus $F^{\prime}\prec F_{j}$ by Proposition 3.5. Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
@@
We have ${F_{q}}^{c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. Recall that, if $F_{q}\in M(\Delta_{3}(W_{n}))$, then $F_{r}=F_{q}$ satisfies ($\ast$3.1). Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $2m-\alpha_{j}\leq\alpha_{i}<\alpha_{j}$ by Remark 3.1 (iv). Further, $2m-\alpha_{j}<m$. This means that either $2m-\alpha_{j}\leq\alpha_{i}\leq m$, or $m<\alpha_{i}<\alpha_{j}$.
@@
In this case also, we show that either $F_{q}\in M(\Delta_{3}(W_{n}))$, or $F_{q}\notin M(\Delta_{3}(W_{n}))$ and we find some $F_{r}\neq F_{q}$ that satisfies ($\ast$3.1).
@@
    Observe that if $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{i}$, then $\alpha_{i}\leq m$ implies that $j_{1}\nsim\alpha_{i}$, and hence $F_{q}\in M(\Delta_{3}(W_{n}))$. So assume that $j_{1}\sim j_{2}$ and $j_{2}\sim\alpha_{i}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let ${F^{\prime}}^{c}:=\{\alpha_{i}\}\sqcup\{j_{1},\alpha_{j}\}$. We first show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\sim j_{2}$ implies that $j_{1}\not\sim\alpha_{j}$. Therefore, if $\alpha_{i}\nsim\alpha_{j}$ or $\alpha_{i}\nsim j_{1}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $\alpha_{i}\sim\alpha_{j}$. We show that $\alpha_{i}\nsim j_{1}$. Suppose $\alpha_{i}\sim j_{1}$. Since $\alpha_{i}\leq m$ and $\alpha_{j}>m$, $\alpha_{i}\sim\alpha_{j}$ implies that $\alpha_{j}\in\{m+1,m+2\}$. If $\alpha_{j}=m+1$, then $\alpha_{i}\in\{m-1,m\}$. Using the facts that $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}<\alpha_{j}$, it follows that $j_{2}\leq m-2$. Hence $j_{1}<m-2$. Then $\alpha_{i}=m-1$, $j_{1}=m-3=\alpha_{j}-4$ and $j_{2}=m-2=\alpha_{j}-3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{3})$, a contradiction.
    If $\alpha_{j}=m+2$, then $\alpha_{i}=m$ and $j_{1}<j_{2}\leq m-3$ (as $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}<\alpha_{j}$). Thus $\alpha_{i}\nsim j_{1}$, which contradicts our assumption that $\alpha_{i}\sim j_{1}$. Since we get a contradiction in each case, our assumption that $\alpha_{i}\sim j_{1}$ is false. Therefore, $\alpha_{i}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$.

### v2

@@ 3.1. Shelling Order
    Now, let $F^{\prime\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, i.e., $C_{n}^{2}[F^{\prime\prime c}]$ is connected. We have $i_{1}<j_{1}<\alpha_{j}$ and $22\leq\alpha_{j}\leq n-$ 2. Suppose $j_{1}\nsim i_{1}$ or $j_{1}\nsim\alpha_{j}$. Then $i_{1}\sim\alpha_{j}$ such that $i_{1}=0$ and $\alpha_{j}=n-2$. Therefore, $\alpha_{j}<_{\mathcal{O}}i_{2}$ implies that $i_{2}\in\{1,2,n-1\}$ by definition of $\mathcal{O}$. We get a contradiction to the fact that $C_{n}^{2}[F_{i}^{c}]$ is disconnected.
@@
    Thus, $j_{1}\sim i_{1}$ and $j_{1}\sim\alpha_{j}$. Now, since $C_{n}^{2}[F_{j}^{c}]$ is disconnected, $j_{2}\nsim\alpha_{j}$, which implies that $j_{2}\geq\alpha_{j}+3$. We have ${F^{\prime}}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Let $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $i_{1}<j_{1}<\alpha_{j}$ and $j_{2}\geq\alpha_{j}+3$, $F^{\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, and thus $F^{\prime}\prec F_{j}$ by Definition 3.5 (i). Hence, by taking $F_{r}=F^{\prime}$, ($\ast$3.1) is satisfied.
@@
    We now assume that $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, i.e., $C_{n}^{2}[F^{\prime c}]$ is connected. Since $j_{2}\nsim\alpha_{j}$, we get $j_{2}\sim i_{1}$ and $i_{1}\sim\alpha_{j}$. Then $i_{1}<j_{1}<\alpha_{j}<j_{2}$ implies that $i_{1}=\alpha_{j}-2$ and $j_{1}=\alpha_{j}-1$. Also, $j_{2}\sim i_{1}$ implies that either $j_{2}=n-2$ and $i_{1}=0$, or $j_{2}=n-1$ and $i_{1}\in\{0,1\}$. If $i_{1}=0$, then $\alpha_{j}=2$, and thus, $\alpha_{j}<_{\mathcal{O}}i_{2}$ implies that $i_{2}\in\{1,n-2,n-1\}$ by definition of $\mathcal{O}$. This means that $C_{n}^{2}[F_{i}^{c}]$ is connected, a contradiction. So, $i_{1}=1$, $j_{1}=2$, $\alpha_{j}=3$ and $j_{2}=n-1$. Further, $C_{n}^{2}[F_{i}^{c}]$ is disconnected implies that $i_{2}\nsim\alpha_{j}$, and since $i_{2}\neq j_{2}$, $5<i_{2}<n-1=j_{2}$. It follows that $i_{2}\nsim j_{1}$. We have ${F^{\prime\prime\prime}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Clearly, $C_{n}^{2}[F^{\prime\prime\prime c}]\in M(\Delta_{3}(C_{n}^{2}))$. Since $j_{1}=\alpha_{j}-1$, $i_{2}\geq\alpha_{j}+3$ and $\alpha_{j}=3\leq m-2$ (as $m\geq 5$), we have $F^{\prime\prime\prime}\notin\mathcal{D}$. Also, $i_{2}<j_{2}$ implies that $F^{\prime\prime\prime}\ll F_{j}$, and thus $F^{\prime\prime\prime}\prec F_{j}$ by Definition 3.5 (i). Hence, $F_{r}=F^{\prime\prime\prime}$ satisfies ($\ast$3.1).
@@
We have ${F_{i}}^{c}=\{\alpha_{i}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$ such that $F_{i}\prec F_{j}$. Then $F_{j}\notin\mathcal{D}$ and $F_{i}\notin\mathcal{T}_{j_{0}}$ implies that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ by Definition 3.4 (ii) and Definition 3.5. Therefore, since $\alpha_{j}<_{\mathcal{O}}j_{1},j_{2}$ by Remark 3.2, we get $\alpha_{i}<_{\mathcal{O}}j_{1},j_{2}$.
Let ${F_{q}}^{c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. Observe that if $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$, then $F_{q}\prec F_{j}$ by Proposition 3.6, and hence $F_{r}=F_{q}$ satisfies ($\ast$3.1) by Proposition 3.7.
@@
In this case, we show that $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$, and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1).
@@
    We have $j_{1}<j_{2}<\alpha_{j}<\alpha_{i}<m<n-1$. Thus $j_{1}\nsim\alpha_{i}$. Also, since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$. It follows that $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$.
@@
    Since $\alpha_{j}<\alpha_{i}<m$, $\alpha_{j}\leq m-2$. Further, $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}>\alpha_{j}$ implies that $j_{1}\geq m+2$ by definition of $\mathcal{O}$. We have $\alpha_{j}<\alpha_{i}<m<m+2\leq j_{1}<j_{2}$. Thus $j_{1}\nsim\alpha_{i}$. Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$. Hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$.
@@
    Since $\alpha_{j}<\alpha_{i}<m$, $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}>\alpha_{j}$, we get $j_{2}\geq m+2$. We have $j_{1}<\alpha_{j}<\alpha_{i}<m<m+2\leq j_{2}$. Thus $j_{2}\nsim\alpha_{i}$. Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
In this case, we show that either $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1), or $F_{q}\notin M(\Delta_{3}(C_{n}^{2}))$ and there exists some $F_{r}\neq F_{q}$ that satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
    Observe that since $j_{1},j_{2}<\alpha_{j}$, if $\alpha_{j}<4$, then $C_{n}^{2}[F_{j}^{c}]$ is connected. So $\alpha_{j}\geq 4$, and hence $\alpha_{i}<2m-\alpha_{j}\leq 2m-4\leq n-3$. We have $j_{1}<j_{2}<\alpha_{j}<m\leq\alpha_{i}<n-3$. Thus $j_{1}\nsim\alpha_{i}$. Since $C_{n}^{2}[F_{j}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{1}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$; and if $j_{2}\nsim\alpha_{j}$, then $j_{2}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$. Hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$.
@@
    Observe that if $j_{1}\nsim\alpha_{i}$ or $j_{1}\nsim j_{2}$, then $\alpha_{i}\geq m$ implies that $j_{2}\nsim\alpha_{i}$, and hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$. So assume that $j_{1}\sim\alpha_{i}$ and $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(C_{n}^{2}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{\alpha_{j},j_{2}\}$. We first show that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$.
    Since $C_{n}^{2}[F_{j}^{c}]$ is disconnected, $j_{1}\sim j_{2}$ implies that $j_{1}\nsim\alpha_{j}$ and $j_{2}\nsim\alpha_{j}$. Therefore, if $\alpha_{i}\nsim\alpha_{j}$ or $\alpha_{i}\nsim j_{2}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Let $\alpha_{i}\sim\alpha_{j}$. We need to show that $\alpha_{i}\nsim j_{2}$. Since $\alpha_{j}<m$ and $\alpha_{i}\geq m$, $\alpha_{i}\sim\alpha_{j}$ implies that $\alpha_{j}\in\{m-2,m-1\}$. If $\alpha_{j}=m-2$, then $\alpha_{i}=m$ and $j_{2}>j_{1}\geq m+2$ (as $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}>\alpha_{j}$). Thus $\alpha_{i}\nsim j_{2}$.
@@
    Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, from Proposition 3.6, we see that $F^{\prime}\prec F_{j}$. Thus $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
@@
    Suppose $j_{2}\nsim\alpha_{i}$. Now, since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, we have $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. If $j_{1}\nsim j_{2}$, then $j_{2}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$; and if $j_{1}\nsim\alpha_{j}$, then $j_{1}\nsim\alpha_{i}$, which implies that $\alpha_{i}$ is an isolated vertex in $C_{n}^{2}[F_{q}^{c}]$. Thus, we see that $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$.
    Now, we assume that $j_{2}\sim\alpha_{i}$. Suppose $j_{1}\sim\alpha_{i}$. Then, $j_{1}<\alpha_{j}<m\leq\alpha_{i}<j_{2}$ implies that either $j_{1}=m-2$, $\alpha_{j}=m-1$, $\alpha_{i}=m$ and $j_{2}\in\{m+1,m+2\}$, or $j_{1}=0$, $\alpha_{i}=n-2$, and $j_{2}=n-1$. In the former case, since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{2}=m+2=\alpha_{j}+3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{2})$, a contradiction as $F_{j}\notin\mathcal{D}$. Hence, we have $j_{1}=0$, $\alpha_{i}=n-2$, and $j_{2}=n-1$. Now, $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}<m$ implies that $\alpha_{j}\leq 2$. It follows that $C_{n}^{2}[{F_{j}}^{c}]$ is connected, again a contradiction. Therefore, $j_{1}\nsim\alpha_{i}$. Now, if $j_{1}\nsim j_{2}$, then $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$.
    So, let $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(C_{n}^{2}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{j_{1},\alpha_{j}\}$. Since $C_{n}^{2}[F_{j}^{c}]$ is disconnected and $j_{1}\sim j_{2}$, we get $j_{1}\nsim\alpha_{j}$. Therefore, $j_{1}\nsim\alpha_{i}$ implies that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. We have $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$. Thus $F^{\prime}\prec F_{j}$ by Proposition 3.6. Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
@@
We have ${F_{q}}^{c}=\{\alpha_{i}\}\sqcup\{j_{1},j_{2}\}$. Recall that, if $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$, then $F_{r}=F_{q}$ satisfies ($\ast$3.1). Since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, $2m-\alpha_{j}\leq\alpha_{i}<\alpha_{j}$ by Remark 3.1 (iv). Further, $2m-\alpha_{j}<m$. This means that either $2m-\alpha_{j}\leq\alpha_{i}\leq m$, or $m<\alpha_{i}<\alpha_{j}$.
@@
In this case also, we show that either $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$, or $F_{q}\notin M(\Delta_{3}(C_{n}^{2}))$ and we find some $F_{r}\neq F_{q}$ that satisfies ($\ast$3.1).
@@
    Observe that if $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{i}$, then $\alpha_{i}\leq m$ implies that $j_{1}\nsim\alpha_{i}$, and hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$. So assume that $j_{1}\sim j_{2}$ and $j_{2}\sim\alpha_{i}$. Then $F_{q}\notin M(\Delta_{3}(C_{n}^{2}))$. Let ${F^{\prime}}^{c}:=\{\alpha_{i}\}\sqcup\{j_{1},\alpha_{j}\}$. We first show that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$.
    Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\sim j_{2}$ implies that $j_{1}\not\sim\alpha_{j}$. Therefore, if $\alpha_{i}\nsim\alpha_{j}$ or $\alpha_{i}\nsim j_{1}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Let $\alpha_{i}\sim\alpha_{j}$. We show that $\alpha_{i}\nsim j_{1}$. Suppose $\alpha_{i}\sim j_{1}$. Since $\alpha_{i}\leq m$ and $\alpha_{j}>m$, $\alpha_{i}\sim\alpha_{j}$ implies that $\alpha_{j}\in\{m+1,m+2\}$. If $\alpha_{j}=m+1$, then $\alpha_{i}\in\{m-1,m\}$. Using the facts that $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}<\alpha_{j}$, it follows that $j_{2}\leq m-2$. Hence $j_{1}<m-2$. Then $\alpha_{i}=m-1$, $j_{1}=m-3=\alpha_{j}-4$ and $j_{2}=m-2=\alpha_{j}-3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{3})$, a contradiction.
    If $\alpha_{j}=m+2$, then $\alpha_{i}=m$ and $j_{1}<j_{2}\leq m-3$ (as $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $j_{2}<\alpha_{j}$). Thus $\alpha_{i}\nsim j_{1}$, which contradicts our assumption that $\alpha_{i}\sim j_{1}$. Since we get a contradiction in each case, our assumption that $\alpha_{i}\sim j_{1}$ is false. Therefore, $\alpha_{i}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$.